OpenStudy (anonymous):

Help pleaseeeeee.. Find the area of the rectangle GIVEN: Diagonal 194 Perimeter of the rectangle =548

OpenStudy (anonymous):



OpenStudy (anonymous):

\[Perimeter = 2(l+w)\] \[Area_\rectangle = w \times l\] l = length w=width

OpenStudy (anonymous):